OpenStudy (anonymous):

can any one give me partial derivative help what is the partial derivative of 6xe^3x square

OpenStudy (anonymous):

This is a function of one variable, isn't it?

OpenStudy (anonymous):

not this is 2 variable function whole question is here f(x,y)=6xe^3x square

OpenStudy (anonymous):

4 viewing but no one replying give me please answer...

OpenStudy (wasiqss):

listen isnt that the fact that partial derivative means it will contain both variables in a same eq, bt here it is only one,

OpenStudy (anonymous):

OK. If you want to find the partial derivative with respect to one variable, you consider the other variable as a constant. That's \(\frac{d}{dx}f(x,y)=6x(3e^{3x})+6e^{3x}=6e^{3x}(3x+1)\), and \(\frac{d}{dy}f(x,y)=0\).

OpenStudy (anonymous):

yes i get the first partial derivative and 2nd gone i want to get this number again partial derivative 2nd term is y w.r.t.x it become zero keeping y constant

OpenStudy (anonymous):

I don't understand. What do you mean?

OpenStudy (anonymous):

ok i write again this question whole

OpenStudy (anonymous):

f(x,y)=6xe^3x this could be treated as a single variable value and by using product rule =3.6xe^3x+e^3x.6 =18xe^3x+6e^3x =6e^3x(3x+1)

OpenStudy (anonymous):

e^3x^2 simple tell what is the partial derivative of this ?

OpenStudy (anonymous):

You mean \(\large e^{3x^2}\)?

OpenStudy (anonymous):

yes here is another value attached with this e^y^2 and i want to take it's partial derivative

OpenStudy (anonymous):

You're confusing everybody altafhpatel!

OpenStudy (anonymous):

and the whole question is f(x,y)=

OpenStudy (anonymous):

sorry i m new here

OpenStudy (anonymous):

Just write the whole function \(\large f(x,y)\), and state clearly what you want to find.

OpenStudy (anonymous):

ok i draw

OpenStudy (anonymous):